OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown?

pooja195 (pooja195):

Following....

OpenStudy (luigi0210):

Do you have any choices..?

OpenStudy (anonymous):

7x−3y=4;2x−4y=1 Step: Solve 7x−3y=4 for x: 7x−3y+3y=4+3y(Add 3y to both sides) 7x=3y+4 7x7=3y+47(Divide both sides by 7) x=37y+47 Step: Substitute (37y+47) for x in 2x−4y=1: 2x−4y=1 2(37y+47)−4y=1 −227y+87=1(Simplify both sides of the equation) −227y+87+−87=1+−87(Add (-8)/7 to both sides) −227y=−17 −227y−227=−17−227(Divide both sides by (-22)/7) y=122 Step: Substitute (122) for y in x=37y+47: x=37y+47 x=37(122)+47 x=1322(Simplify both sides of the equation) Answer: x=1322 and y=122